# Bayesian Inference for Consistent Predictions in Overparameterized Nonlinear Regression

### v1

@@ 3. Main Results
In this section, we present the main theoretical results. We begin by introducing the key assumptions that underlie our analysis (§3.1). These assumptions concern the spectral properties of the covariance matrix and the concentration of the empirical covariance matrix around its population counterpart. We then proceed to establish the posterior contraction for the GLM (§3.2). Our theoretical results offer characterizations of the posterior contraction, elucidating the intricate interplay between the dimensionality, sample size, and the eigenvalue decay of the covariance matrix.
@@ 3.1. Assumptions
We begin by focusing on the boundedness of the operator norm of $\Sigma$ and the concentration of the empirical covariance matrix $\widehat{\Sigma}$ around $\Sigma$.
@@
This assumption ensures that the empirical covariance matrix concentrates well around the population covariance matrix. The sequence $\rho_{n}$ quantifies the rate of this concentration, while $\ell_{n}$ controls the probability of the concentration event. It is worth noting that this assumption is mild and includes not only the sub-Gaussian distributions commonly imposed in high-dimensional statistics but also log-concave distributions, which encompass most unimodal parametric distributions [56, 57] and allow for heavy-tailed distributions such as the Laplace distribution. In contrast, other related literature has imposed more restrictive conditions such as sub-Gaussianity [5] or strong log-concavity [13]. In a high-dimensional setting, where many covariates are incorporated, the mildness of the distributional assumption is particularly important, as it allows for a wider range of applicable scenarios.
Our next assumption concerns the behavior of the eigenvalues of $\Sigma$.
@@
Given a covariance matrix $\Sigma$ satisfying (9), with eigenvalues $\{\lambda_{k}\}_{k=1}^{p}$, the sequence $\{\rho_{n}\}_{n\in\mathbb{N}}$ from Assumption 1, and a sequence of bounds $(L_{\kappa},U_{\kappa})$, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ of positive reals such that the following hold as $n\to\infty$:
@@
This assumption imposes certain regularity conditions on the eigenvalues of $\Sigma$. Intuitively, conditions (i) and (ii) control the rate at which the bulk of the eigenvalues (those indexed from $L_{\kappa}$ to $U_{\kappa}$) are allowed to decay moderately. Condition (iii) requires that the eigenvalues beyond the bulk (those indexed larger than $U_{\kappa}$) are sufficiently small, while condition (iv) constrains the size of $U_{\kappa}$ relative to the sample size $n$ and the rate $\varepsilon_{n}$. These conditions are crucial for deriving the posterior contraction rates in Theorem 1.
Finally, we assume a certain local Lipschitz condition of the link function $g$ in terms of the Kullback-Leibler (KL) divergence and the variational form between the corresponding probability distributions.
@@
$K(P,P^{\prime}),~{}V(P,P^{\prime})\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
@@
This assumption limits the amount of variation in a distribution when different parameters are used, meaning that it controls the complexity of the model. This assumption is satisfied by several common GLMs, as illustrated in the following examples.
@@ 3.2. Posterior Contraction
Consider the GLM (2) with covariate distribution satisfying Assumptions 1-3, and a prior distribution on $\bm{\beta}$ is (6). Then, for a sufficiently large constant $M$, the posterior distribution contracts around the true parameter $\bm{\beta}^{*}$ as
[…]
@@
This theorem shows that under suitable conditions on the covariance matrix $\Sigma$ and the prior distribution, the posterior distribution of the predictor concentrates around the oracle predictor at a rate $\varepsilon_{n}$ with respect to the $L_{2}(P_{\bm{X}})$ distance. The rate $\varepsilon_{n}$ depends on the interplay between the sample size $n$, the concentration rate $\rho_{n}$ of the empirical covariance matrix, and the eigenvalues of $\Sigma$ in the bulk region.
@@ 3.2.1. Consistency as a distribution
While Theorem 1 derives posterior consistency of the predictor with respect to the $L_{2}$ distance, it also holds for other commonly used distances between probability distributions, such as the total variation distance, the Hellinger distance, and the KL divergence in several cases, including a list of distributions in Example 1. The contraction in these divergences is particularly useful for assessing the quality of the posterior distribution in terms of its ability to approximate the true data-generating distribution. A detailed discussion of these results is provided in the appendix.
@@ 3.2.2. Consistency of Point Estimators
Under the same assumptions as in Theorem 1, it can be shown that the MAP estimator is consistent with respect to the $L_{2}(P_{\bm{X}})$ norm:
@@
Under the assumptions of Theorem 1 and the additional condition that posterior consistency with respect to the total variation distance or the Hellinger distance holds (as discussed in §3.2.1), the posterior mean is also consistent:
@@
As the posterior mean considers the weights of the entire distribution (as it is defined by integral), its consistency requires the additional condition that the posterior distribution is not ill-behaved. We defer the further discussion to the appendix.
@@ 4. Extension to Two-Parameter Gaussian Case
In this section, we extend our analysis to the setting where the GLM has a Gaussian distribution for the measurement error. Specifically, we consider the following model, referred to as the single neuron model [47, 60]:
@@
Here, we specify the prior distributions for the parameters $\bm{\beta}$ and $\sigma^{2}$. For $\bm{\beta}$, we continue to use the prior distribution (6). Regarding the variance parameter, we employ as the prior distribution the inverse Gaussian distribution $\Pi_{\varsigma^{2}}(\cdot)$ with the density of the form:
@@
where $\eta>0$ and $\xi>0$ denote the mean and shape parameters, respectively. We allow for arbitrary assignment of hyperparameters. An important property of the inverse Gaussian prior is that the tails on both sides decrease rapidly, which is critical for posterior contraction. In contrast, the inverse gamma distribution, the conjugate prior to the Gaussian likelihood, has a light lower tail and a heavy upper tail. From a Bayesian perspective, the inverse Gaussian prior reflects the analyst’s belief that the parameter will not have extremely large or small values. This prior distribution for the variance parameter has also been employed in [51, 41, 59].
@@
Consider the regression model (16) with covariate distribution satisfying Assumptions 1-3, and a prior distribution on $(\bm{\beta},\sigma^{2})$ satisfying (6) and (17). Then, for a sufficiently large constant $M$, the following convergences hold as $n\to\infty$:
@@
This theorem provides the posterior contraction for both the predictor $g(\bm{X}^{\top}\bm{\beta})$ and the error variance $\sigma^{2}$ in the two-parameter Gaussian model. The contraction rate for the predictor is with respect to the $L_{2}(P_{\bm{X}})$ distance, while the rate for $\sigma^{2}$ is with respect to the absolute difference.
The extension to the two-parameter Gaussian model is particularly relevant for applications where the error variance is unknown and needs to be estimated from the data. By considering the joint posterior contraction rates for $(\bm{\beta},\sigma^{2})$, Theorem 2 presents a theoretical justification for Bayesian inference in such models and offers guidance on the choice of prior distributions that can adapt to the unknown error variance.
@@ 5. Simulation
This section confirms by simulation the statement of the theorem presented in §3 and §4.
@@ 5.1. Logistic Regression
Here, we consider a binary classification scenario where the response variable $Y_{i}$ takes values in $\{0,1\}$. The response variables are generated according to the following model:
$Y_{i}=\mathrm{Ber}(g_{ls}(\bm{X}_{i}^{\top}\bm{\beta}));\quad i=1,2,...,n,$
@@
where $g_{ls}(r):=1/(1+e^{-r})$ is the logistic function. We set $\bm{\beta}\sim N(\bm{0},I_{p})$ and $p=n^{4/3}$, ensuring an overparameterized setting. Regarding the covariate, we consider two settings: (A) Gaussian covariate, that is, $\bm{X}_{i}\sim N(\bm{0},\Sigma)$; (B) Laplacian convariate, that is, $\bm{X}_{i}\sim L(\bm{0},\Sigma)$, where the eigenvalues of $\Sigma$ follow $\lambda_{j}(\Sigma)=10\exp(-j/8)+n\exp\{-\sqrt{n}\}/p$.
@@
We investigate the behavior of the proposed method as the sample size increases by varying the training data size $n$ from $50$ to $500$. For each sample size, we generate a test dataset of $1000$ instances to evaluate the performance of the trained classifier. To evaluate the performance of the Bayesian predictions, we employ four widely used metrics: 0-1 loss, which quantifies the misclassification rate; Area Under the ROC Curve (AUC), which assesses the discriminative ability of the classifier; Unconfident Misclassifications rate (UM), which measures the proportion of unconfident predictions among misclassified instances; and Confident Correct predictions rate (CC), which calculates the proportion of accurate predictions among confident ones. We refer to a confident prediction as one in which the $95\%$ interval does not cross the threshold $(0.5)$, while an unconfident one does. The latter two metrics are helpful when building systems that defer action when there is no confidence in the prediction.
@@
Figure 1. Change of 0-1 loss, AUC (area under the ROC curve), UM (proportion of unconfident predictions among misclassifications), and CC (proportion of accurate predictions among confident ones) as the sample size grows in the logistic regression with the Gaussian covariate case (left) and the Laplace covariate case (right).
@@
Figure 1 illustrates the results of the experiment. As the sample size increases, the 0-1 loss exhibits a consistent decrease in both settings, indicating an improvement in classification accuracy. This trend persists even in the overparameterized setting, aligning with the theoretical findings. The AUC metric also shows a steady increase with the sample size, suggesting an enhanced ability to discriminate between the two classes. The UM and CC metrics provide insights into the uncertainty estimation of the proposed method. UM consistently presents a high value. This indicates that the method is able to warn us of a lack of confidence when misclassifying. The CC also presents a similarly high value, which means that confident predictions are trustworthy. These observations suggest that the proposed method effectively captures the uncertainty associated with predictions and leverages it to make more informed decisions.
@@ 5.2. Nonlinear regression with Gaussian noise
In this subsection, we consider a scenario where the true response takes values in the positive real line $\mathbb{R}_{+}$. Specifically, the response variables are generated as follows:
$Y_{i}=N(g_{sp}(\bm{X}_{i}^{\top}\bm{\beta}),1);\quad i=1,2,...,n,$
where $\bm{\beta}\sim N(\bm{0},I_{p})$, $p=n^{4/3}$, and $g_{sp}(r):=\log(1+e^{r})$ is the softplus function. The distributions of the covariate vectors are set to follow the same patterns as in the previous experiment.
We examine how the performance of Bayesian predictions changes as the data size increases by varying the training data size $n$ from $50$ to $500$. For a newly generated test dataset $\{(\bm{x}_{i},g_{sp}(\bm{x}_{i}^{\top}\bm{\beta}))\}_{i=1}^{1000}$, we assess the accuracy of Bayesian predictions using three widely adopted indicators: Root Mean Squared Error (RMSE), measuring the average deviation between the predicted values and the true values; Coverage Probability (CP), evaluating the reliability of the $95\%$ prediction intervals by calculating the proportion of true values falling within the intervals; and Average Length (AL), quantifying the width of the $95\%$ prediction intervals, providing insight into the precision of the predictions.
Figure 2. Change of RMSE (root mean squared error), CP (coverage probability of $95$% prediction interval), and AL (average length of $95$% prediction interval) as the sample size grows in nonlinear Gaussian regression with the Gaussian covariate case (left) and the Laplace covariate case (right).
The results are shown in Figure 2. First, the RMSE decreases monotonically as the sample size increases. Even when $p$ becomes large and the setting becomes clearly overparameterized, the fact that the accuracy continues to improve supports the theoretical results experimentally. Next, we focus on CP and AL. These are indicators related to prediction intervals, with CP representing the reliability of the prediction interval and AL representing its width. From the figure, CP is not good when $n=100$, but after that, it stabilizes at a high value. AL is also slightly increasing, but not too much, implying that the interval is not trivial. (Regardless of $n$, the mean of $\{g_{sp}(\bm{x}_{i}^{\top}\bm{\beta})\}_{i=1}^{1000}$ in the test data is around $3.5$ and the variance is around $5$.) These results suggest that the proposed method is able to properly quantify the uncertainty in constructing prediction intervals.

### v2

@@ 3. Main Results
In this section, the main theoretical results are presented. Initially, the key assumptions that underlie our analysis (§3.1) are introduced. These assumptions concern the spectral properties of the covariance matrix and the concentration of the empirical covariance matrix around its population counterpart. Subsequently, the posterior contraction for the GLM is established (§3.2). The theoretical results characterize the posterior contraction, elucidating the intricate interplay between dimensionality, sample size, and eigenvalue decay of the covariance matrix.
@@ 3.1. Assumptions
The initial focus is on the boundedness of the operator norm of $\Sigma$ and concentration of the empirical covariance matrix $\widehat{\Sigma}$ around $\Sigma$.
@@
This assumption ensures that the empirical covariance matrix concentrates around the population covariance matrix. The sequence $\rho_{n}$ quantifies the rate of this concentration, whereas $\ell_{n}$ controls the probability of the concentration event. Notably, this assumption is mild and includes not only the sub-Gaussian distributions commonly imposed in high-dimensional statistics but also log-concave distributions, which encompass most unimodal parametric distributions [57, 58] that allow for heavy-tailed distributions such as the Laplace distribution. In contrast, other related literature impose more restrictive conditions such as sub-Gaussianity [6] or strong log-concavity [14]. In a high-dimensional setting, where many covariates are incorporated, the mildness of the distributional assumption is particularly important, as it allows for a wider range of applicable scenarios.
The next assumption concerns the behavior of the eigenvalues of $\Sigma$.
@@
Given a covariance matrix $\Sigma$ satisfying (9), with eigenvalues $\{\lambda_{k}\}_{k=1}^{p}$, sequence $\{\rho_{n}\}_{n\in\mathbb{N}}$ from Assumption 1, and a sequence of bounds $(L_{\kappa},U_{\kappa})$, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ of positive reals such that the following hold as $n\to\infty$:
@@
This assumption imposes certain regularity conditions on the eigenvalues of $\Sigma$. Intuitively, conditions (i) and (ii) control the rate at which the bulk of the eigenvalues (those indexed from $L_{\kappa}$ to $U_{\kappa}$) are allowed to decay moderately. Condition (iii) requires that the eigenvalues beyond the bulk (those indexed larger than $U_{\kappa}$) are sufficiently small, whereas condition (iv) constrains the size of $U_{\kappa}$ relative to sample size $n$ and rate $\varepsilon_{n}$. These conditions are crucial for deriving the posterior contraction rates in Theorem 1.
Finally, we assume a certain local Lipschitz condition of the link function $g$ in terms of the Kullback-Leibler (KL) divergence and variational form between the corresponding probability distributions.
@@
$K(P,P^{\prime}),V(P,P^{\prime})\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
@@
This assumption limits the amount of variation in a distribution when different parameters are used, to control the complexity of the model and is satisfied by several common GLMs, as illustrated in the following examples.
@@ 3.2. Posterior Contraction
Consider the GLM (2) with covariate distribution satisfying Assumptions 1-3, and the prior distribution on $\bm{\beta}$ is (6). Then, for a sufficiently large constant $M$, the posterior distribution contracts around the true parameter $\bm{\beta}^{*}$ as
$\Pi\Big{(}{\bm{\beta}:\|g(\bm{X}^{\top}\bm{\beta})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}>M\varepsilon_{n}}\mid,\mathcal{D}\Big{)}\overset{P^{*}}{\longrightarrow}0\quad\mbox{as}\quad n\to\infty.$
@@
This theorem shows that under conditions suitable for the covariance matrix $\Sigma$ and prior distribution, the posterior distribution of the predictor concentrates around the oracle predictor at a rate $\varepsilon_{n}$ with respect to the $L_{2}(P_{\bm{X}})$ distance. The rate $\varepsilon_{n}$ depends on the interplay between sample size $n$, concentration rate $\rho_{n}$ of the empirical covariance matrix, and eigenvalues of $\Sigma$ in the bulk region.
@@ 3.2.1. Consistency as a distribution
While Theorem 1 derives the posterior consistency of the predictor with respect to the $L_{2}$ distance, it also holds for other commonly used distances between probability distributions, such as the total variation distance, the Hellinger distance, and the KL divergence in several cases, including a list of distributions in Example 1. The contraction in these divergences is particularly useful for assessing the quality of the posterior distribution in approximating the true data-generating distribution.
@@ 3.2.2. Consistency of Point Estimators
Under the same assumptions as in Theorem 1, the MAP estimator can be shown to be consistent with respect to the $L_{2}(P_{\bm{X}})$ norm:
@@
Under the assumptions of Theorem 1 and the additional condition that posterior consistency with respect to the total variation distance or Hellinger distance holds (as discussed in §3.2.1), the posterior mean is also consistent:
@@
As the posterior mean considers the weights of the entire distribution (defined by integral), its consistency requires the additional condition that the posterior distribution is not ill-behaved. Further discussion is deferred to the Appendix.
@@ 4. Extension to Two-Parameter Gaussian Case
In this section, the analysis is extended to the setting of GLM with a Gaussian distribution for the measurement error. Specifically, we consider the single-neuron model [48, 61]:
@@
Here, after specifying the prior distributions for the parameters $\bm{\beta}$ and $\sigma^{2}$, the prior distribution (6) for $\bm{\beta}$ is used. Regarding the variance parameter, we employ as the prior distribution, the inverse Gaussian distribution $\Pi_{\varsigma^{2}}(\cdot)$ with density of the form:
@@
where $\eta>0$ and $\xi>0$ denote the mean and shape parameters, respectively, and the hyperparameters are arbitrarily assigned. An important property of the inverse Gaussian prior is that the tails on both sides decrease rapidly, which is critical for posterior contraction. In contrast, the inverse gamma distribution, the conjugate prior to the Gaussian likelihood, has a light lower tail and a heavy upper tail. From a Bayesian perspective, the inverse Gaussian prior reflects the analyst’s belief that the parameter will not have extremely large or small values. This prior distribution for the variance parameter has also been employed in [52, 42, 60].
@@
Consider the regression model (15) with covariate distribution satisfying Assumptions 1-3 and a prior distribution on $(\bm{\beta},\sigma^{2})$ satisfying (6) and (16). Then, for a sufficiently large constant $M$, the following convergences hold as $n\to\infty$:
@@
This theorem provides the posterior contraction for both predictor $g(\bm{X}^{\top}\bm{\beta})$ and error variance $\sigma^{2}$ in the two-parameter Gaussian model. The contraction of the predictor is with respect to the $L_{2}(P_{\bm{X}})$ distance, while that of $\sigma^{2}$ is with respect to the absolute difference.
The extension of the two-parameter Gaussian model is particularly relevant to applications where the error variance is unknown and needs to be estimated from the data. By considering the joint posterior contraction rates for $(\bm{\beta},\sigma^{2})$, Theorem 2 presents a theoretical justification for Bayesian inference in such models and offers guidance on the choice of prior distributions that can adapt to the unknown error variance.
@@ 5. Simulation
This section confirms by simulation the statement of the theorem presented in §3 and §4 through simulations. The posterior predictive distribution of the proposed method is computed using variational inference, and the detailed algorithm is deferred to the Appendix.
@@ 5.1. Logistic Regression
Here, we consider a binary classification scenario where the response variable $Y_{i}$ takes values in $\{0,1\}$. The data are generated according to the following model:
$Y_{i}=\mathrm{Ber}(g_{ls}(\bm{X}_{i}^{\top}\bm{\beta}));\quad i=1,2,...,n,$ (20)
@@
where $g_{ls}(r):=1/(1+e^{-r})$ is the logistic function. Specifically, for $n=50,100,\ldots,500$, to ensure an overparameterized setting, $\bm{\beta}\sim N(\bm{0},I_{p})$ are generated with $p=n^{4/3}$. Two settings are considered for the covariate: (A) Gaussian covariate, $\bm{X}_{i}\sim N(\bm{0},\Sigma)$; (B) Laplacian covariate, $\bm{X}_{i}\sim L(\bm{0},\Sigma)$, where the eigenvalues of $\Sigma$ follow $\lambda_{j}(\Sigma)=10\exp(-j/8)+n\exp\{-\sqrt{n}\}/p$. Then, each $Y_{i}$ corresponding to $X_{i}$ and $\beta$ is sampled from (20).
@@
For every $n$, $20$ different datasets were repeatedly generated through the above procedure and each was analyzed to determine the behavior of the proposed method in response to an increase in sample size. In each case, a test dataset of $1000$ instances was generated, to evaluate the performance of the trained classifier. The performance of the Bayesian predictions was evaluated, employing four widely used metrics: 0-1 loss, which quantifies the misclassification rate; area under the ROC curve (AUC), which assesses the discriminative ability of the classifier; rate of unconfident misclassifications (UM), which measures the proportion of unconfident predictions among misclassified instances; and confident correct (CC) predictions rate, which calculates the proportion of accurate predictions among confident ones. We refer to a confident prediction as one whose $95\%$ predictive interval does not cross the threshold $(0.5)$, whereas an unconfident one does. The latter two metrics are helpful when building systems that defer action when there is no confidence in the prediction. Additionally, principal component regression (PCR) [5] was performed as a baseline method, with the number of principal components selected by 5-fold cross-validation from $\{1,2,\ldots,30\}$.
@@
Figure 1. Change of 0-1 loss, AUC (area under the ROC curve), UM (proportion of unconfident predictions among misclassifications), and CC (proportion of accurate predictions among confident ones) of the proposed method as sample size increases for the cases of logistic regression with Gaussian covariate (left) and Laplace covariate (right). The gray lines represent the results of the principal component regression. The shaded bands represent the 90% interval.
@@
Figure 1 illustrates the results, where the reported values are averaged over the $20$ datasets. As sample size increases, the 0-1 loss exhibits a consistent decrease in both settings, indicating an improvement in classification accuracy. This trend persists even in the overparameterized setting, aligning with the theoretical findings. The AUC metric also shows a steady increase with sample size, suggesting an enhanced ability to discriminate between the two classes. Our method outperformed PCR, and these results underscore the importance of overparameterization for prediction. The UM and CC metrics provide insights into the uncertainty estimation of the proposed method. UM consistently presents a high value. This indicates that the method can warn us of a lack of confidence when misclassifying. The CC also presents a similarly high value, suggesting that confident predictions are trustworthy. These observations suggest that the proposed method effectively captures the uncertainty associated with predictions, thereby leveraging it to make more informed decisions.
@@ 5.2. Nonlinear regression with Gaussian noise
In this subsection, the scenario of true responses taking values along the positive real line $\mathbb{R}_{+}$ is considered. Specifically, the data are generated as follows:
$Y_{i}=N(g_{sp}(\bm{X}_{i}^{\top}\bm{\beta}),1);\quad i=1,2,...,n,$ (21)
where $g_{sp}(r):=\log(1+e^{r})$ is the softplus function. For $n=50,100,\ldots,500$, we set $p=n^{4/3}$ and sample the coefficient $\bm{\beta}\in\mathbb{R}^{p}$ and covariate vectors $X_{i}\in\mathbb{R}^{p}$ in the same manner as in the previous experiment. Then, the scalar response is generated by (21).
For every $n$, $20$ distinct datasets were generated using the above procedure and, on each one, Bayesian prediction performance changes were examined with an increase in training data size. For a newly generated test dataset $\{(\bm{x}_{i},g_{sp}(\bm{x}_{i}^{\top}\bm{\beta}))\}_{i=1}^{1000}$, the accuracy of Bayesian predictions was assessed using three widely adopted indicators: root-mean-squared error (RMSE), measuring the average deviation between predicted and true values; coverage probability (CP), evaluating the reliability of the $95\%$ prediction intervals by calculating the proportion of true values falling within the confidence intervals; and average length (AL), quantifying the width of the $95\%$ prediction intervals, providing insight into the precision of the predictions. Moreover, PCR was implemented, selecting the number of principal components using 5-fold cross-validation from $\{1,2,\ldots,30\}$.
Figure 2. Change of RMSE (root mean squared error), CP (coverage probability of $95$% prediction interval), and AL (average length of $95$% prediction interval) as sample size increases in nonlinear Gaussian regression with the Gaussian covariate case (left) and Laplace covariate case (right).
The results, averaged over $20$ datasets, are shown in Figure 2. First, the RMSE decreases monotonically as sample size increases. Even when $p$ is large and the setting is clearly overparameterized, the fact that accuracy continues to improve supports the theoretical results experimentally. Also, our method excels PCR, providing advantages of our method for prediction problems. Next, we focus on CP and AL, which are indicators related to prediction intervals, with CP representing the reliability of the prediction interval and AL representing its width. In the figure, the coverage is not sufficient when $n=100$; however, after that, CP stabilizes at a high value. AL also slightly increases but not by much, implying that the interval is not trivial. (Regardless of $n$, the mean of $\{g_{sp}(\bm{x}_{i}^{\top}\bm{\beta})\}_{i=1}^{1000}$ in the test data is around $3.5$, and the variance is around $5$.) These results suggest that the proposed method properly quantifies the uncertainty in constructing prediction intervals.